# Deformations of pseudocharacters and Mazur’s finiteness condition

### v1

@@ 3. Geometric invariant theory
We first recall the set up of [15]. Let $\mathcal{O}$ be a quasi-excellent ring and let $S=\operatorname{Spec}\mathcal{O}$. Then $\mathcal{O}$ is Nagata [16, Tag 07QV] and hence universally Japanese [16, Tag 0334]. The last assumption is imposed in [15]. Let $G$ be a reductive group scheme over $S$, so that $G$ is an affine group scheme over $S$, $G\rightarrow S$ is smooth and the geometric fibres are connected reductive groups.
Let $V$ be a free $\mathcal{O}$-module of finite rank $r$ endowed with a $G$-module structure, let $\check{V}=\operatorname{Hom}_{\mathcal{O}}(V,\mathcal{O})$ and let $\operatorname{Sym}(\check{V})$ be the symmetric algebra over $\mathcal{O}$. The $G$-module structure on $V$ induces an action of $G$ on Spec($\operatorname{Spec}(\operatorname{Sym}(\check{V}))=\mathbb{A}^{r}_{S}.$ Let $X$ be a closed $G$-invariant subscheme of $\operatorname{Spec}(\operatorname{Sym}(\check{V}))$. The $G$-action on $X$ induces an action on $B$, the ring of functions on $X$. The GIT quotient $X\sslash G$ is represented by the ring of invariants $B^{G}$. Below we let $R$ be a $B^{G}$-algebra, which is a quasi-excellent ring. In the application, both $\mathcal{O}$ and $R$ will be complete noetherian local rings, and these are quasi-excellent [16, Tag 07QW].
@@
If $R$ is a flat $B^{G}$-algebra then [15, Lemma 2] implies that
@@
and we are done. If $R$ is flat over $\mathcal{O}$ then [15, Lemma 2] implies that
@@
It follows from [15, Theorem 2 (ii)] that $(R\otimes_{B^{G}}B)^{G}$ is a finite $(R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V}))^{G}$-module. Since the image of $\operatorname{Sym}(\check{V})^{G}$ in $B$ is contained in $B^{G}$, the action of $R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V})^{G}$ on $(R\otimes_{B^{G}}B)^{G}$ factors through the map $R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V})^{G}\rightarrow R\otimes_{\mathcal{O}}B^{G}\rightarrow R$. Hence, $(R\otimes_{B^{G}}B)^{G}$ is a finite $R$-module. ∎
@@
3.1 implies $(R\otimes_{B^{G}}B)^{G}$ is a finite $R$-module. Since $R$ is a local noetherian ring this implies that every closed subscheme of $(X\times_{X\sslash G}\operatorname{Spec}R)\sslash G$ contains a point above the closed point of $\operatorname{Spec}R$. Since $G$ is connected every irreducible component of $Z$ is $G$-invariant by [3, Lemma 2.1]. Thus we may assume that $Z$ is irreducible. The image of $Z$ in $(X\times_{X\sslash G}\operatorname{Spec}R)\sslash G$ is closed by [15, Theorem 3 (iii)] and thus $Z$ contains a point $x$ above the closed point of $\operatorname{Spec}R$. Since $X\rightarrow X\sslash G$ is of finite type we may choose $x$ such that $\kappa(x)$ is finite over $\kappa$. ∎
@@
Let $\mathfrak{p}$ be the prime ideal of $B^{G}$ corresponding to $y$ and let $R$ be the completion of the localisation $(B^{G})_{\mathfrak{p}}$. We note that $R$ is complete local noetherian and hence quasi-excellent and also $B^{G}$-flat. Since $B$ is normal, $B^{G}$ and $(B^{G})_{\mathfrak{p}}$ are normal. Since $\mathcal{O}$ is quasi-excellent and $B^{G}$ is of finite type over $\mathcal{O}$ by [15, Theorem 2], $(B^{G})_{\mathfrak{p}}$ is quasi-excellent [16, Tag 07QU]. Since $(B^{G})_{\mathfrak{p}}$ is quasi-excellent and normal, its completion $R$ is also normal [16, Tag 0C23], and hence an integral domain.
Let $Q$ be the fraction field of $R$. Since $R$ is $B^{G}$-flat and $Q$ is $R$-flat, $R$ is a subring of $B_{R}:=R\otimes_{B^{G}}B$ and $Q$ is a subring of $B_{Q}:=Q\otimes_{B^{G}}B$. In particular, $X_{R}:=\operatorname{Spec}B_{R}$ and $X_{Q}:=\operatorname{Spec}B_{Q}$ are non-empty. Let $Z^{\prime}$ be an irreducible component of $X_{Q}$ corresponding to a prime ideal $\mathfrak{q}^{\prime}$ of $B_{Q}$. Then $Z^{\prime}$ is $G$-invariant by [3, Lemma 2.1]. Let $Z$ be the closure of $Z^{\prime}$ in $X_{R}$. Then $Z$ is a closed $G$-invariant subset of $X_{R}$, corresponding to the prime ideal $\mathfrak{q}^{\prime\prime}=\mathfrak{q}^{\prime}\cap B_{R}$.
3.2 implies that there is $\bar{x}_{1}\in Z(\Omega)$ above $y$. Let $\bar{x}_{2}\in Z(\Omega)$ be a point contained in the closure of the $G$-orbit of $\bar{x}_{1}$ in $Z\otimes_{R}\Omega$, such that its $G$-orbit has minimal dimension. Then $\bar{x}_{2}\in Z(\Omega)$ lies above $y$ and its $G$-orbit is closed. It follows from [15, Theorem 3 (ii)] that $\bar{x}$ and $\bar{x}_{2}$ lie in the same $G$-orbit and thus $\bar{x}\in Z(\Omega)$.
@@ 4. Finitely generated free profinite groups
Let $F$ be a free group on $r$ generators $\gamma_{1},\ldots,\gamma_{r}$. Let $L$ be a finite extension with the ring of integers $\mathcal{O}$ and residue field $k$. Let $G$ be a generalised reductive group scheme over $\mathcal{O}$.
@@
By [14, Theorem 3.20], $\mathrm{PC}^{F}_{G}$ is representable by $\mathop{\mathstrut\rm colim}\limits\nolimits_{F_{m}\to F}\mathcal{O}[G^{m}]^{G^{0}}$, where the colimit is indexed by the category of finitely generated free groups $F_{m}$ together with a homomorphism $F_{m}\to F$ and homomorphisms of free groups commuting with the homomorphisms to $F$. This category has a terminal object, given by an isomorphism $F_{r}\xrightarrow{\cong}F$. It follows, that the colimit is isomorphic to $\mathcal{O}[G^{r}]^{G^{0}}$. ∎
Let $\overline{\Theta}\in\mathrm{PC}^{F}_{G}(k)$ and let $\overline{\rho}:F\rightarrow G(\bar{k})$ be a $G$-semisimple representation such that $\Theta_{\overline{\rho}}=\overline{\Theta}\otimes_{k}\bar{k}$, such representation exists and is uniquely determined up to $G^{0}(\bar{k})$-conjugation by [14, Theorem 3.7].
@@
The map $\rho\mapsto(\rho(\gamma_{1}),\ldots,\rho(\gamma_{r}))$ induces an isomorphism of $\mathcal{O}$-schemes $\operatorname{Rep}^{F}_{G}\cong G^{r}$. Since $G$ is smooth over $\mathcal{O}$, we deduce that $\operatorname{Rep}^{F}_{G}$ is smooth over $\mathcal{O}$. Since $\mathcal{O}$ is a DVR we deduce that $\operatorname{Rep}^{F}_{G}$ is regular and hence normal. Since $\overline{\rho}$ is $G$-semisimple the $G$-orbit of $\bar{x}\in\operatorname{Rep}^{F}_{G}(\bar{k})$ corresponding to $\overline{\rho}$ is closed, [12, Proposition 6.7]. The assertion follows from 4.1 and 3.3 applied with $X=\operatorname{Rep}^{F}_{G}$. ∎
@@
Similarly $\operatorname{Def}^{F}_{\overline{\Theta}}$ is pro-represented by $\mathcal{O}(\mathrm{PC}^{F}_{G})_{\mathfrak{m}}^{\wedge}$. Since $F$ is dense in $\widehat{F}$ the restriction $\Theta\mapsto\Theta|_{F}$ induces an injection $\operatorname{Def}^{\widehat{F}}_{\overline{\Theta}}(A)\hookrightarrow\operatorname{Def}^{F}_{\overline{\Theta}}(A)$ for all $A\in\mathfrak{A}_{\mathcal{O}}$ by [14, Lemma 3.2]. This yields a surjection $\mathcal{O}(\mathrm{PC}^{F}_{G})_{\mathfrak{m}}^{\wedge}\twoheadrightarrow R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}$.
@@
If $A\in\mathfrak{A}_{\mathcal{O}}$ then the kernel of $G(A)\rightarrow G(k)$ is a finite $p$-group. Hence, if $\rho_{A}$ is a deformation of $\overline{\rho}$ to $A$ then the homomomorhim $\rho_{A}:\widehat{F}\rightarrow G(A)$ factors through $Q\rightarrow G(A)$. We deduce that the surjection $R^{\square,\widehat{F}}_{\overline{\rho}}\twoheadrightarrow R^{\square,Q}_{\overline{\rho}}$ is an isomorphism. Since the natural map $R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}\rightarrow R^{\square,\widehat{F}}_{\overline{\rho}}$ is injective by 4.3, we deduce that the composition
@@ 5. The general case
Let $\Gamma$ be a profinite group and let $\overline{\Theta}\in\mathrm{cPC}^{\Gamma}_{G}(k)$. Let $R^{\mathrm{ps}}:=R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ be the universal ring representing $\operatorname{Def}_{\overline{\Theta}}$, and let $\Theta^{u}$ be the universal deformation of $\overline{\Theta}$. By the reconstruction theorem [14, Theorem 3.7] there exists a $G$-semisimple representation $\overline{\rho}:\Gamma\rightarrow G(\bar{k})$, uniquely determined up to $G^{0}(\bar{k})$-conjugation, such that $\Theta_{\overline{\rho}}=\overline{\Theta}\otimes_{k}\bar{k}$. Theorem 3.8 in [14] implies that $\overline{\rho}$ is continuous. We let $K$ be the kernel of $\overline{\rho}$. Since the topology on $\bar{k}$ is discrete $K$ is an open subgroup of $\Gamma$. Since $\overline{\rho}$ is uniquely determined up to $G^{0}(\bar{k})$-conjugation $K$ is independent of the choice of $\overline{\rho}$. Moreover, we have inclusions $\operatorname{Ker}(\Theta^{u})\subset\operatorname{Ker}(\overline{\Theta})\subset K$ and $\operatorname{Ker}(\Theta^{u})$ is a closed normal subgroup of $\Gamma$.
@@
Let $N$ be a closed normal subgroup of $\Gamma$ contained in $K$. If for every finite subset $Y\subset X$ the map $K_{Y}\rightarrow\Gamma/N$ factors through the maximal pro-$p$ quotient $K_{Y}(p)$ then $K/N$ is pro-$p$. In particular, $\Gamma\twoheadrightarrow\Gamma/N$ factors through $Q\twoheadrightarrow\Gamma/N$.
@@ 6. Mazur’s finiteness condition
we deduce that $\operatorname{Hom}^{\mathrm{cont}}(K(p),\mathbb{F}_{p})$ is a finite dimensional $\mathbb{F}_{p}$-vector space. This implies that $K(p)$ is finitely generated, [10, Lemma 2.1]. The assertion follows as $\Gamma/K$ is a finite group. ∎
@@
The assertion follows from the corresponding result for finitely generated profinite groups [14, Theorem 5.7]. ∎
@@ 6.1. The moduli space of condensed representations
Let $R$ be a profinite noetherian $\mathcal{O}$-algebra. Then any finitely generated $R$-module $M$ carries a unique Hausdorff topology, which makes it into a topological $R$-module. Let $A$ be an $R$-algebra. A representation $\rho:\Gamma\rightarrow G(A)$ is called $R$-condensed if there exists a closed immersion of $\mathcal{O}$-schemes $\tau:G\hookrightarrow\mathbb{A}^{n}$ and a finitely generated $R$-submodule $M$ of $\mathbb{A}^{n}(A)=A^{n}$ such that $\tau(\rho(\Gamma))\subset M$ and the map $\tau\circ\rho:\Gamma\rightarrow M$ is continuous for the canonical topology on $M$ as a finitely generated $R$-module. We check in [12, Lemma 4.2] that if $\rho$ is $R$-condensed with respect to one closed immersion $\tau$ then the same holds for all closed immersions. Moreover, in [12, Section 4.2] we show that $\rho$ is $R$-condensed if and only if there exists a homomorphism of condensed groups $\underline{\Gamma}\to G(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})$ which recovers $\rho$ after evaluation at a point. Here $\underline{-}$ denotes the functor from topological spaces to condensed sets and $(-)_{\operatorname{disc}}$ takes a set to a discrete condensed set. This interpretation in terms of condensed mathematics of Clausen and Scholze motivates the terminology. Informally, $R$-condensed should be thought of as a continuity condition on $\rho$, even if there is no topology on the $R$-algebra $A$.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then $R^{\mathrm{ps}}_{\overline{\Theta}}$ is noetherian by 6.2 and for an $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra $A$ we let $X^{\mathrm{gen}}_{\overline{\Theta}}(A)$ be the set of $R^{\mathrm{ps}}_{\overline{\Theta}}$-condensed representations $\rho:\Gamma\rightarrow G(A)$ such that $\Theta_{\rho}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}A$.
@@
We choose a closed immersion of $\mathcal{O}$-group schemes $\tau:G\hookrightarrow\operatorname{GL}_{d}$. Composition with $\tau$ induces injections $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}(A)\hookrightarrow X^{\mathrm{gen},\Gamma}_{\tau\circ\overline{\Theta}}(A)$ and $X^{\mathrm{gen},Q}_{\overline{\Theta}}(A)\hookrightarrow X^{\mathrm{gen},Q}_{\tau\circ\overline{\Theta}}(A)$ by [12, Proposition 3.8]. Thus we may assume that $G=\operatorname{GL}_{d}$. In this case the assertion follows from [12, Lemma 5.5]. ∎

### v2

@@ 3. Geometric invariant theory
We first recall the set up of [18]. Let $\mathcal{O}$ be a quasi-excellent ring and let $S=\operatorname{Spec}\mathcal{O}$. Then $\mathcal{O}$ is Nagata [19, Tag 07QV] and hence universally Japanese [19, Tag 0334]. The last assumption is imposed in [18]. Let $G$ be a reductive group scheme over $S$, so that $G$ is an affine group scheme over $S$, $G\rightarrow S$ is smooth and the geometric fibres are connected reductive groups.
Let $V$ be a free $\mathcal{O}$-module of finite rank $r$ endowed with a $G$-module structure, let $\check{V}=\operatorname{Hom}_{\mathcal{O}}(V,\mathcal{O})$ and let $\operatorname{Sym}(\check{V})$ be the symmetric algebra over $\mathcal{O}$. The $G$-module structure on $V$ induces an action of $G$ on $\operatorname{Spec}(\operatorname{Sym}(\check{V}))=\mathbb{A}^{r}_{S}.$ Let $X$ be a closed $G$-invariant subscheme of $\operatorname{Spec}(\operatorname{Sym}(\check{V}))$. The $G$-action on $X$ induces an action on $B$, the ring of functions on $X$. The GIT quotient $X\sslash G$ is represented by the ring of invariants $B^{G}$. Below we let $R$ be a $B^{G}$-algebra, which we assume to be a quasi-excellent ring. In the application, both $\mathcal{O}$ and $R$ will be complete noetherian local rings, and these are quasi-excellent [19, Tag 07QW].
@@
If $R$ is a flat $B^{G}$-algebra then [18, Lemma 2] implies that
@@
and we are done. If $R$ is flat over $\mathcal{O}$ then [18, Lemma 2] implies that
@@
It follows from [18, Theorem 2 (ii)] that $(R\otimes_{B^{G}}B)^{G}$ is a finite $(R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V}))^{G}$-module. Since the image of $\operatorname{Sym}(\check{V})^{G}$ in $B$ is contained in $B^{G}$, the action of $R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V})^{G}$ on $(R\otimes_{B^{G}}B)^{G}$ factors through the map $R\otimes_{\mathcal{O}}\operatorname{Sym}(\check{V})^{G}\rightarrow R\otimes_{\mathcal{O}}B^{G}\rightarrow R$. Hence, $(R\otimes_{B^{G}}B)^{G}$ is a finite $R$-module. ∎
@@
3.1 implies $(R\otimes_{B^{G}}B)^{G}$ is a finite $R$-module. Since $R$ is a local noetherian ring this implies that every closed subscheme of $(X\times_{X\sslash G}\operatorname{Spec}R)\sslash G$ contains a point above the closed point of $\operatorname{Spec}R$. Since $G$ is connected every irreducible component of $Z$ is $G$-invariant by [4, Lemma 2.1]. Thus we may assume that $Z$ is irreducible. The image of $Z$ in $(X\times_{X\sslash G}\operatorname{Spec}R)\sslash G$ is closed by [18, Theorem 3 (iii)] and thus $Z$ contains a point $x$ above the closed point of $\operatorname{Spec}R$. Since $X\rightarrow X\sslash G$ is of finite type we may choose $x$ such that $\kappa(x)$ is finite over $\kappa$. ∎
@@
Let $\mathfrak{p}$ be the prime ideal of $B^{G}$ corresponding to $y$ and let $R$ be the completion of the localisation $(B^{G})_{\mathfrak{p}}$. We note that $R$ is complete local noetherian and hence quasi-excellent and also $B^{G}$-flat. Since $B$ is normal, $B^{G}$ and $(B^{G})_{\mathfrak{p}}$ are normal. Since $\mathcal{O}$ is quasi-excellent and $B^{G}$ is of finite type over $\mathcal{O}$ by [18, Theorem 2], $(B^{G})_{\mathfrak{p}}$ is quasi-excellent [19, Tag 07QU]. Since $(B^{G})_{\mathfrak{p}}$ is quasi-excellent and normal, its completion $R$ is also normal [19, Tag 0C23], and hence an integral domain.
Let $Q$ be the fraction field of $R$. Since $R$ is $B^{G}$-flat and $Q$ is $R$-flat, $R$ is a subring of $B_{R}:=R\otimes_{B^{G}}B$ and $Q$ is a subring of $B_{Q}:=Q\otimes_{B^{G}}B$. In particular, $X_{R}:=\operatorname{Spec}B_{R}$ and $X_{Q}:=\operatorname{Spec}B_{Q}$ are non-empty. Let $Z^{\prime}$ be an irreducible component of $X_{Q}$ corresponding to a prime ideal $\mathfrak{q}^{\prime}$ of $B_{Q}$. Then $Z^{\prime}$ is $G$-invariant by [4, Lemma 2.1]. Let $Z$ be the closure of $Z^{\prime}$ in $X_{R}$. Then $Z$ is a closed $G$-invariant subset of $X_{R}$, corresponding to the prime ideal $\mathfrak{q}^{\prime\prime}=\mathfrak{q}^{\prime}\cap B_{R}$.
3.2 implies that there is $\bar{x}_{1}\in Z(\Omega)$ above $y$. Let $\bar{x}_{2}\in Z(\Omega)$ be a point contained in the closure of the $G$-orbit of $\bar{x}_{1}$ in $Z\otimes_{R}\Omega$, such that its $G$-orbit has minimal dimension. Then $\bar{x}_{2}\in Z(\Omega)$ lies above $y$ and its $G$-orbit is closed. It follows from [18, Theorem 3 (ii)] that $\bar{x}$ and $\bar{x}_{2}$ lie in the same $G$-orbit and thus $\bar{x}\in Z(\Omega)$.
@@ 4. Finitely generated free profinite groups
Let $F$ be a free group on $r$ generators $\gamma_{1},\ldots,\gamma_{r}$. Let $L$ be a finite extension of $\mathbb{Q}_{p}$ with the ring of integers $\mathcal{O}$ and residue field $k$. Let $G$ be a generalised reductive group scheme over $\mathcal{O}$.
@@
By [16, Theorem 3.20], $\mathrm{PC}^{F}_{G}$ is representable by $\mathop{\mathstrut\rm colim}\limits\nolimits_{F_{m}\to F}\mathcal{O}[G^{m}]^{G^{0}}$, where the colimit is indexed by the category of finitely generated free groups $F_{m}$ together with a homomorphism $F_{m}\to F$ and homomorphisms of free groups commuting with the homomorphisms to $F$. This category has a terminal object, given by an isomorphism $F_{r}\xrightarrow{\cong}F$. It follows, that the colimit is isomorphic to $\mathcal{O}[G^{r}]^{G^{0}}$. ∎
Let $\overline{\Theta}\in\mathrm{PC}^{F}_{G}(k)$ and let $\overline{\rho}:F\rightarrow G(\bar{k})$ be a $G$-semisimple representation such that $\Theta_{\overline{\rho}}=\overline{\Theta}\otimes_{k}\bar{k}$, such representation exists and is uniquely determined up to $G^{0}(\bar{k})$-conjugation by [16, Theorem 3.7].
@@
The map $\rho\mapsto(\rho(\gamma_{1}),\ldots,\rho(\gamma_{r}))$ induces an isomorphism of $\mathcal{O}$-schemes $\operatorname{Rep}^{F}_{G}\cong G^{r}$. Since $G$ is smooth over $\mathcal{O}$, we deduce that $\operatorname{Rep}^{F}_{G}$ is smooth over $\mathcal{O}$. Since $\mathcal{O}$ is a DVR we deduce that $\operatorname{Rep}^{F}_{G}$ is regular and hence normal. Since $\overline{\rho}$ is $G$-semisimple the $G$-orbit of $\bar{x}\in\operatorname{Rep}^{F}_{G}(\bar{k})$ corresponding to $\overline{\rho}$ is closed, [14, Proposition 6.7]. The assertion follows from 4.1 and 3.3 applied with $X=\operatorname{Rep}^{F}_{G}$. ∎
@@
Similarly $\operatorname{Def}^{F}_{\overline{\Theta}}$ is pro-represented by $\mathcal{O}(\mathrm{PC}^{F}_{G})_{\mathfrak{m}}^{\wedge}$. Since $F$ is dense in $\widehat{F}$ the restriction $\Theta\mapsto\Theta|_{F}$ induces an injection $\operatorname{Def}^{\widehat{F}}_{\overline{\Theta}}(A)\hookrightarrow\operatorname{Def}^{F}_{\overline{\Theta}}(A)$ for all $A\in\mathfrak{A}_{\mathcal{O}}$ by [16, Lemma 3.2]. This yields a surjection $\mathcal{O}(\mathrm{PC}^{F}_{G})_{\mathfrak{m}}^{\wedge}\twoheadrightarrow R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}$.
@@
If $A\in\mathfrak{A}_{\mathcal{O}}$ then the kernel of $G(A)\rightarrow G(k)$ is a finite $p$-group. Hence, if $\rho_{A}$ is a deformation of $\overline{\rho}$ to $A$ then the homomorphism $\rho_{A}:\widehat{F}\rightarrow G(A)$ factors through $Q\rightarrow G(A)$. We deduce that the surjection $R^{\square,\widehat{F}}_{\overline{\rho}}\twoheadrightarrow R^{\square,Q}_{\overline{\rho}}$ is an isomorphism. Since the natural map $R^{\mathrm{ps},\widehat{F}}_{\overline{\Theta}}\rightarrow R^{\square,\widehat{F}}_{\overline{\rho}}$ is injective by 4.3, we deduce that the composition
@@ 5. The general case
Let $\Gamma$ be a profinite group and let $\overline{\Theta}\in\mathrm{cPC}^{\Gamma}_{G}(k)$. Let $R^{\mathrm{ps}}:=R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ be the universal ring representing $\operatorname{Def}_{\overline{\Theta}}$, and let $\Theta^{u}$ be the universal deformation of $\overline{\Theta}$ introduced in Section 2.3. By the reconstruction theorem [16, Theorem 3.7] there exists a $G$-semisimple representation $\overline{\rho}:\Gamma\rightarrow G(\bar{k})$, uniquely determined up to $G^{0}(\bar{k})$-conjugation, such that $\Theta_{\overline{\rho}}=\overline{\Theta}\otimes_{k}\bar{k}$. Theorem 3.8 in [16] implies that $\overline{\rho}$ is continuous. We let $K$ be the kernel of $\overline{\rho}$. Since the topology on $\bar{k}$ is discrete, $K$ is an open subgroup of $\Gamma$. Since $\overline{\rho}$ is uniquely determined up to $G^{0}(\bar{k})$-conjugation, $K$ is independent of the choice of $\overline{\rho}$. Moreover, we have inclusions $\operatorname{Ker}(\Theta^{u})\subset\operatorname{Ker}(\overline{\Theta})\subset K$ and $\operatorname{Ker}(\Theta^{u})$ is a closed normal subgroup of $\Gamma$.
@@
Let $N$ be a closed normal subgroup of $\Gamma$ contained in $K$. If for every finite subset $Y\subset X$ the map $K_{Y}\rightarrow\Gamma/N$ factors through the maximal pro-$p$ quotient $K_{Y}(p)$, then $K/N$ is pro-$p$. In particular, $\Gamma\twoheadrightarrow\Gamma/N$ factors through $Q\twoheadrightarrow\Gamma/N$.
@@ 6. Mazur’s finiteness condition
we deduce that $\operatorname{Hom}^{\mathrm{cont}}(K(p),\mathbb{F}_{p})$ is a finite dimensional $\mathbb{F}_{p}$-vector space. This implies that $K(p)$ is finitely generated, [11, Lemma 2.1]. The assertion follows as $\Gamma/K$ is a finite group. ∎
@@
The assertion follows from the corresponding result for finitely generated profinite groups [16, Theorem 5.7]. ∎
@@ 6.1. The moduli space of condensed representations
Let $R$ be a profinite noetherian $\mathcal{O}$-algebra. Then any finitely generated $R$-module $M$ carries a unique Hausdorff topology, which makes it into a topological $R$-module. Let $A$ be an $R$-algebra. A representation $\rho:\Gamma\rightarrow G(A)$ is called $R$-condensed if there exists a closed immersion of $\mathcal{O}$-schemes $\tau:G\hookrightarrow\mathbb{A}^{n}$ and a finitely generated $R$-submodule $M$ of $\mathbb{A}^{n}(A)=A^{n}$ such that $\tau(\rho(\Gamma))\subset M$ and the map $\tau\circ\rho:\Gamma\rightarrow M$ is continuous for the canonical topology on $M$ as a finitely generated $R$-module. We check in [14, Lemma 4.2] that if $\rho$ is $R$-condensed with respect to one closed immersion $\tau$ then the same holds for all closed immersions. Moreover, in [14, Section 4.2] we show that $\rho$ is $R$-condensed if and only if there exists a homomorphism of condensed groups $\underline{\Gamma}\to G(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})$ which recovers $\rho$ after evaluation at a point. Here $\underline{-}$ denotes the functor from topological spaces to condensed sets and $(-)_{\operatorname{disc}}$ takes a set to a discrete condensed set. This interpretation in terms of condensed mathematics of Clausen and Scholze [17] motivates the terminology. Informally, $R$-condensed should be thought of as a continuity condition on $\rho$, even if there is no topology on the $R$-algebra AA.
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then $R^{\mathrm{ps}}_{\overline{\Theta}}$ is noetherian by 6.2. Thus we may apply the discussion above with $R=R^{\mathrm{ps}}_{\overline{\Theta}}$ and for an $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra $A$ we let $X^{\mathrm{gen}}_{\overline{\Theta}}(A)$ be the set of $R^{\mathrm{ps}}_{\overline{\Theta}}$-condensed representations $\rho:\Gamma\rightarrow G(A)$ such that $\Theta_{\rho}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}A$.
@@
We choose a closed immersion of $\mathcal{O}$-group schemes $\tau:G\hookrightarrow\operatorname{GL}_{d}$. Composition with $\tau$ induces injections $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}(A)\hookrightarrow X^{\mathrm{gen},\Gamma}_{\tau\circ\overline{\Theta}}(A)$ and $X^{\mathrm{gen},Q}_{\overline{\Theta}}(A)\hookrightarrow X^{\mathrm{gen},Q}_{\tau\circ\overline{\Theta}}(A)$ by [14, Proposition 3.8]. Thus we may assume that $G=\operatorname{GL}_{d}$. In this case the assertion follows from [14, Lemma 5.5]. ∎